OpenStudy (anonymous):

Scores on a certain exam are normally distributed with a mean of 78 and a variance of 25. What is the z-value for a score of 90?

OpenStudy (anonymous):

(Standard Deviation)^2=Variance=25 that means the Standard Deviation = 5 so Z=(x-mean)/Standard Deviation this is the formula you have to memorize it so Z=(90-78)/5 = 12/5 = 2.4 if you need any more help with Statistics i will be glad to help

OpenStudy (anonymous):

thanks, I appreciate it. I need help with this other problem to...

OpenStudy (anonymous):

http://www.psychassessment.com.au/pdf/mod2.pdf Read the above for more insights.

OpenStudy (anonymous):

yes sure just post it here :)

OpenStudy (anonymous):

thanks eliassaab.

OpenStudy (anonymous):

The average length of time between arrivals at a turnpike tollbooth is 23 seconds. Assume that the time between arrivals at the tollbooth is exponentially distributed. (a) What is the probability that a minute or more will elapse between arrivals? (b) If a car has just passed through the tollbooth, what is the probability that no car will show up for at least 3 minutes?

OpenStudy (anonymous):

ohh :S exponentially distributed ! i have no idea whats that maybe i forgot it! or we didnt take it yet sorry

OpenStudy (anonymous):

it's all good. yeah, I'm taking business stats. right now.

OpenStudy (anonymous):

ohh i took the engineering one so maybe there is a difference or something